The velocity of a river of wide 400m is 2 km per hour a Boat is rowed from a point on the bank to exactly opposite point on the other bank with a velocity of a 8kilometre per hour distance between the exactly opposite point and the point which the bank us reaches the bank is

Grade:8

1 Answers

Arun
5 years ago
Here the boat moves with the resultant velocity R. But the vertical component 8 km/h takes him to the opposite shore. Tan θ = 2/8 = 1/4 Velocity = 8 km/h distance = 0.4 km Time = 0.4/8 = 0.05 sec. The boat will reach at point C. In ∆ABC, tan θ = BC/AB = BC/0.4 = 1/4 ⇒ BC = 0.1 km= 100 m.